OpenStudy (anonymous):

A CYLINDER OF COMPRESSED GAS HAS A VOLUME OF 60L AND A PRESSURE OF 10 ATM. WHAT VOLUME WOULD GAS OCCUPY IF ALLOWED TO ESCAPE INTO A BALLOON AT A PRESSURE OF 1.5 ATM?

OpenStudy (anonymous):

i KNW I HAVE TO CONVERT BUT IDK WAT TO CONVERT FIRST

OpenStudy (anonymous):

@Ashleyisakitty

OpenStudy (anonymous):

@AnImEfReaK

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

ITS BOYLES AND DALTONS LAW

OpenStudy (anonymous):

V1P1=V2P2- use this formula. You change atm to kpA. 1 atm = 101.3 kpa. You leave volume in litres as it is.

OpenStudy (anonymous):

OH I FORGOT TO MENTION ASSUME THERE IS NO TEMPERATURE CHANGE

OpenStudy (anonymous):

WOULD THE V1 BE THE ORIGINAL VOLUME?

OpenStudy (anonymous):

Ok no problem. If 1 atm = 101.3 Kpa then... 10 atm * 101.3 kpa / 1atm. That's your first conversion of 10 atm into kpa. (21) then... 1.5 atm * 101.3 kpa/1atm . That's your second convertion of 10 atm into kpa. (P2)

OpenStudy (anonymous):

V1 is the volume they gave you. You are figuring out V2.

OpenStudy (anonymous):

OK HOLD ON LET ME WRITE THIS DOWN

OpenStudy (anonymous):

Alright :)

OpenStudy (anonymous):

P1** NOT 21 typing error ***

OpenStudy (anonymous):

OKAY SO FOR P1 I GOT 1,013 KPA

OpenStudy (anonymous):

AND P2 I GOT 151.95 KPA

OpenStudy (anonymous):

V1 IS 60 L

OpenStudy (anonymous):

HOW DO U FIND V2?

OpenStudy (anonymous):

Now find V2 by substituting those known values into the equation.

OpenStudy (anonymous):

V1*P1=V2*P2 just re-arrange to solve for V2.

OpenStudy (anonymous):

SO I MULTIPLY THE FIRST TWO NUMBERS AND DIVIDE BY THE SECOND #

OpenStudy (anonymous):

Can you type that up for me in equation format?

OpenStudy (anonymous):

Like how I did above ^

OpenStudy (anonymous):

SO 60*1,013=60,780/151.95=400 L =V2

OpenStudy (anonymous):

Yes, that is correct :)

OpenStudy (anonymous):

OH OK HEY CAN U HELP WITH ONE MORE ITS CAUSE THEY ASK ME TO USE A CHART AND IDK HOW TO USE IT AND THANKS:)

OpenStudy (anonymous):

Alright ^^

OpenStudy (anonymous):

THIS IS THE CHART

OpenStudy (anonymous):

So what's the question?

OpenStudy (anonymous):

THIS IS THE QUESTION

OpenStudy (anonymous):

#9

OpenStudy (anonymous):

Alright did you get the first right?

OpenStudy (anonymous):

YEA

OpenStudy (anonymous):

Then what are you having difficulty with exactly? If you could do that question the rest will be easy.

OpenStudy (anonymous):

I FORGOT HOW I SOLVED IT

OpenStudy (anonymous):

Oh, ok. Are you familiar with PV = nRT?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

I HAVENT DONE CHEM IN A WHILE

OpenStudy (anonymous):

Ok P is our pressure, V volume, n moles, R gas constant 8.314 and T temperature is specific.

OpenStudy (anonymous):

Step 1: In the question they gave you T. But its in degrees Celcius. It must be converted to K, Kelvin. 0 degrees Celcius = 273.15 degrees Kelvin so.. 28 degrees Celsius = how many K?

OpenStudy (anonymous):

103.15 KELVIN

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

OH WAIT DONT U SUBTRACT 273 AND 103.15

OpenStudy (anonymous):

Where is the 103.15 coming from?

OpenStudy (anonymous):

IDK I DID MY MATH WRONG BUT I REMEBER MY TEACHER SAID SOMETHING ABOUT SUBTRACTING

OpenStudy (anonymous):

Don't get Kpa confused with Kelvin, use the rule above- do you get it?

OpenStudy (anonymous):

There is addition here 28 + 273.15...

OpenStudy (anonymous):

OH YEA

OpenStudy (anonymous):

Remember 273.15 represents 0 degrees celcius.

OpenStudy (anonymous):

301.15?

OpenStudy (anonymous):

And we are adding 28 to it.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

SO P IN THE FORMULA IS 301.15?

OpenStudy (anonymous):

Step 2: Covert 0.985 atm to Kpa like i taught you in the previous question.

OpenStudy (anonymous):

No that's T temperature.

OpenStudy (anonymous):

That's why you may have got confused...

OpenStudy (anonymous):

OH OK

OpenStudy (anonymous):

Step 2: That will give you P for pressure.

OpenStudy (anonymous):

Do you get it so far?

OpenStudy (anonymous):

400.95 IS P?

OpenStudy (anonymous):

Its 0.985 atm * ? what's your calculation?

OpenStudy (anonymous):

to convert to kPa.

OpenStudy (anonymous):

*101.32500KPA

OpenStudy (anonymous):

1 ATM =101.32500

OpenStudy (anonymous):

So try multiplying again.

OpenStudy (anonymous):

AND I GOT 0.985*101.32500 AND THE CALCULATOR GAVE ME 400.95?

OpenStudy (anonymous):

NVR MIND I GOT 99.80

OpenStudy (anonymous):

Now that's right.

OpenStudy (anonymous):

I got 99.8 on both computer and hand held calculators just to make sure :)

OpenStudy (anonymous):

OK HOW DO U FIND R AND N?

OpenStudy (anonymous):

R is a fixed gas constant. 8.314.

OpenStudy (anonymous):

n is what you are solving for because they gave V.

OpenStudy (anonymous):

So step 3: Convert 624cm^3 into L, if 1L = 1000cm^3

OpenStudy (anonymous):

You devide the 624 by ? to get 1L?

OpenStudy (anonymous):

divide*

OpenStudy (anonymous):

I DIVIDE 1000CM^3 AND 624?

OpenStudy (anonymous):

No you only divide 624 by 1000 here.

OpenStudy (anonymous):

IM CONFUSED SO WAT TO NUMBERS DO I DIVIDE?

OpenStudy (anonymous):

FOR THE PROBLEM

OpenStudy (anonymous):

Okay, If 1L = 1000 CM^3 (We want Volume in Litres here) We are given 624cm^3 to make that into litres, divide it by 1000. See, 624 cm^3 * 1L/ 1000 cm^3.

OpenStudy (anonymous):

Get it so far?

OpenStudy (anonymous):

YES COUNTINUE

OpenStudy (anonymous):

So what is your V in Litres now?

OpenStudy (anonymous):

242,970.624 BUT IT MIGHT BE TRICKY CAUSE ITS AN IPHONE CALCULATOR

OpenStudy (anonymous):

624/ 1000 = 0.624.

OpenStudy (anonymous):

OH ITS CAUSE I INCLUDED THE ^3

OpenStudy (anonymous):

That's just the unit- apart of cm^3.

OpenStudy (anonymous):

OH OK

OpenStudy (anonymous):

You only use the numbers in your calculations.

OpenStudy (anonymous):

Step 4: You have R- its a fixed gas constant- its value is 8.314.

OpenStudy (anonymous):

Now can you list for me the values of V, R, P and T we collected? You need them to find n.

OpenStudy (anonymous):

SO WAIT MY V IS 0.624

OpenStudy (anonymous):

AND I KNW P IS 99.80 T IS 301.15 R IS 8.314

OpenStudy (anonymous):

Okay, substitute then all in the ideal gas law to find n PV = nRT.

OpenStudy (anonymous):

WATS THE UNIT FOR THE R? AND PV=NRT SO (99.80)(0.624L)=N(8.314 )(301.15 KPA)

OpenStudy (anonymous):

OH AND FOR THE FIRST PART I GOT 62.27?

OpenStudy (anonymous):

J/K/mol is the unit of the gas constant.

OpenStudy (anonymous):

WOULD THE FIRST PART BE 62.27 L OR KPA?